OpenStudy (anonymous):

Hello guys, please help me this one f(x)= -x^7 - 3x^3 - t + 5 , find (f^(-1))'(0) Thanks

OpenStudy (anonymous):

Im confused what are you asking whats the" ' " after the brackets mean and why is f^(-1) ? do you mean f(-1) sorry maybe I am of no help here.

OpenStudy (precal):

" ' " denotes derivative, this is a calculus question

OpenStudy (anonymous):

yep, " ' " means derivative

OpenStudy (anonymous):

oh I see you what does the zero in brackets mean?

OpenStudy (anonymous):

and what does f^(-1) I might be able to help if understand what you are asking

OpenStudy (precal):

slope at that point, if you are studying calculus this is standard notation. f^(-1) inverse function

myininaya (myininaya):

\[(f^{-1})'(a)=\frac{1}{f'(f^{-1}(a))}\]

OpenStudy (anonymous):

I'm not bad at it I just thought inverse was f(x)^-1

OpenStudy (anonymous):

is that s'posed to be a "t" in your equation?

myininaya (myininaya):

\[f(x)=-x^7-3x^3-x+5\] We need to find \[f^{-1}(0)\] So what can we plug in that will give us 0?

OpenStudy (anonymous):

Wait are we supposed to take the derivative of the inverse and then solve for zero? is that the question?

OpenStudy (precal):

yes, but you meant evaluate at zero

OpenStudy (anonymous):

lol yeah my bad

OpenStudy (precal):

That is ok. Many problems are posted incorrectly. People write solve instead of evaluate. Some people incorrectly write the powers eg 3x2 instead of 3x^2..

myininaya (myininaya):

well we want to figure out what we can plug into that polynomial that will give us 0 so if we plug in 1 what do we get ? :)

myininaya (myininaya):

so therefore what is \[f^{-1}(0)=?\]

myininaya (myininaya):

\[\text{ Since } f(1)=0 \text{ then } f^{-1}(0)=?\]

OpenStudy (anonymous):

they got cut off myin..

myininaya (myininaya):

what?

myininaya (myininaya):

Oh you mean the user who asked the question isn't here anymore? :(

OpenStudy (anonymous):

you were cut off too for a little while...

OpenStudy (anonymous):

keep going please I am interested

myininaya (myininaya):

What is f inverse of 0 then ?

OpenStudy (anonymous):

1

myininaya (myininaya):

ok\[(f^{-1})'(0)=\frac{1}{f'(f^{-1}(0))}=\frac{1}{f'(1)}\] The last thing to do is find f'(x) Then plug in 1 for x and voo-lah or is it woo-lah ?

myininaya (myininaya):

and put 1 on top of that of course :)

myininaya (myininaya):

of the f'(1)

OpenStudy (anonymous):

ahh makes perfect sense now